# Staying on Top of SMEFT-Likelihood Analyses

### v1

@@ 3 Public likelihoods
Normalization (’normfactor’) $\kappa_{sb}=\mu_{b}$
Table 4: List of modifiers in the construction of the HistFactory likelihoods, adapted from Ref. [145]. Here $g_{p}$ and $f_{p}$ describe different interpolation strategies required to compute the modifiers for any $\alpha$ from the values $\kappa_{sb,\alpha=\pm 1},\Delta_{sb,\alpha=\pm 1}$ provided in the likelihood.
For a standard SFitter analysis we extract systematic uncertainties for each measurement from the respective experimental publications. Systematics of the same type are fully correlated between measurements of the same experiment. This approach has drawbacks, for instance that we can only use the uncertainty categories reported in the experimental publications or on HEPData, and that this information often needs to be extracted by hand. Public likelihoods include the full information on a large number of systematic uncertainties in a documented manner, making their implementation more accurate and efficient.
Likelihoods are published in the HistFactory format [56], similar to the SFitter likelihood in Eq.(8). For each bin $b$ measured in a kinematic distribution of a given channel or final state it provides
@@
with the nominal expected rate $\bar{m}_{sb}$ and multiplicative $(\kappa_{sb})$ and additive $(\Delta_{sb})$ modifiers for each physics process $s$. These modifiers correspond to the nuisance parameters affecting the event rate $m_{b}$. The type of modifier and the constraints on its corresponding nuisance parameter depends on the type of uncertainty, the most common are given in Tab. 4. Using the public likelihoods in terms of modifiers and nominal rates $\bar{m}_{sb}$, we can reproduce the experimental results. For visualization we use cabinetry [146], a python library making use of pyhf for statistical analyses.
Starting from the public likelihoods, we organize the full set of nuisance parameters corresponding to systematic uncertainties in a small number of categories. This allows for an easier numerical treatment at, essentially, no cost. To compute the ranges of nuisance parameters for these categories, we first use a profile likelihood to determine the central values and, in a second step, an analysis of the distribution of the nuisance parameter. In this section we will show how we implement and test three public ATLAS likelihoods.
@@ 3.1 ATLAS $t\bar{t}$ likelihood
To test our implementation and evaluation of the public likelihood, we first reproduce some key ATLAS results in Fig. 2. We show the values for each nuisance parameter that maximizes the likelihood, and the pulls,
@@
Here, $\hat{\theta}$ describes the maximum likelihood values and $\theta_{0}$ the value before the fit, normalized to the pre-fit uncertainty $\Delta\theta$. We also show the impact of the individual nuisance parameters on the signal strength $\mu$. It is determined by repeating the fit after fixing the nuisance parameter to its maximum-likelihood value $\hat{\theta}$, shifted by its prefit (postfit) uncertainties ±$\pm\Delta\theta(\pm\Delta\hat{\theta})$. The left panel of Fig. 2 is taken from Ref. [47], while the right panel shows our reproduced results. Both sets show excellent agreement, with negligible differences for a few select nuisance parameters.
Next, we analyze the full likelihood as a function of a single nuisance parameter. This allows us to check the validity of a Gaussian likelihood, as assumed for systematic uncertainties in SFitter. For each nuisance parameter we generally find excellent agreement with the Gaussian assumption, as seen on the left, with only a few exceptions. Figure 3 shows two such cases, one with excellent agreement and one with poor agreement. Even the larger deviations are under control, showing good agreement with the Gaussian approximation when we translate them into one standard deviation. Our combination of nuisance parameters into categories washes out non-Gaussian shapes in these exceptions.
Finally, we test the correlations between individual nuisance parameters and between nuisance parameters assigned to the categories implemented in SFitter. The left panel of Fig. 4 shows the correlations of all individual nuisance parameters with at least one correlation greater than 0.4. We find that out of the many nuisance parameters included in the public likelihood, only very few are significantly correlated. We mainly see strong correlations, for instance, for modeling choices or jets.
In the standard SFitter approach we group these individual nuisance parameters into uncorrelated categories and implement these categories with a single nuisance parameter each. Standard categories cover leptons, jets, tagging, and luminosity. Additional categories are process-specific, such as certain background or missing transverse energy. For all processes in our dataset we use 21 nuisance parameters describing the systematic uncertainties, assumed to be uncorrelated. Using the full likelihood we show the correlations between these categories in the right panel of Fig. 4. The fact that the correlations between categories essentially vanish, validates this SFitter approach.
@@ 3.2 ATLAS $t\bar{t}Z$ likelihood
The second likelihood we implement is for the $t\bar{t}Z$ rate measurement [48]. It simultaneously fits both 3-lepton and 4-lepton signal regions and the corresponding control regions. The parameter of interest is the $t\bar{t}Z$ signal strength. A total of 230 nuisance parameters describe the systematic uncertainties. Unlike for the $t\bar{t}$ likelihood, there are no uncertainties on the shape of the signal, since each signal region is described by a single bin.
Following the method described for the $t\bar{t}$ analysis, we also test the $t\bar{t}Z$ likelihood and our implementation. Figure 5 compares the impact and pulls taken from Ref. [48] with those reproduced by us. We see excellent agreement for all nuisance parameters, with at most very minor differences.
As before, we then show the correlations between nuisance parameters with at least one correlation greater than 0.3 in the left panel of Fig. 6. We can compare them to our SFitter implementation on the right. We find that the correlations between the individual nuisance parameters are already much smaller for this likelihood. The only strong correlations appear between scale uncertainties and the signal strength of the corresponding background. Consequently, the results, after combining all nuisance parameters into the SFitter categories, display negligible correlations between categories.
@@ 3.3 ATLAS $s$-channel single top likelihood
The correlations in the left panel of Fig. 8 show strong correlations only between select nuisance parameters. The strongest correlations appear between jet-related uncertainties and the signal strengths of the two dominant backgrounds, $t\bar{t}$ and $W$+jets. For SFitter, these nuisance parameters are put into the background uncertainty category. These strong correlations are therefore implicitly included in this larger category and the final implementation into SFitter is essentially uncorrelated, as one can see in the right of Fig. 8. While one still finds nonzero correlation between the jet and background uncertainties, and between the jet and lepton uncertainties, these are all negligibly small.
@@ 4 Global analysis
Using, for the first time, public likelihoods in a global SMEFT analysis, allows us to look at different relevant questions. From the data included in our analysis we know that our global analysis is somewhat unlikely to uncover a fundamental and statistically significant break-down of the SM. We first look at the impact on the constraining power from new measurements, especially boosted top kinematics, relative to Ref. [22]. We then study the impact of correlated uncertainties, encoded in the public likelihoods. From a pure statistics perspective we also check if lower-dimensional limits extracted by profiling and by marginalization differ.
@@ 4.1 Better and boosted measurements
Before we study more conceptual questions of global SMEFT analyses, we update our dataset with new measurements, as marked in Tabs. 2 and 3. In Fig. 9 we show the constraints on a selection of 2-dimensional correlations of Wilson coefficients, using all top data compared to the previous SFitter top analysis [22]. All constraints are the result of an analysis of all 22 Wilson coefficients. To extract limits on pairs of coefficients, we use a profile likelihood. Potential differences to a marginalization will be discussed in Sec. 4.3.
@@
The right panel shows the improvement in constraints on $\mathcal{O}_{\phi Q}^{3}$ and $\mathcal{O}_{Qq}^{31}$. Single top production provides constraints on them, and we again benefit from the public likelihood [49], a new $p_{T,t}$ distribution in $t$-channel single top production [115], and new measurements of the $tW$ production cross section [121, 125]. We observe an improvement of the individual constraints and of their correlation. In particular, $\mathcal{O}_{Qq}^{31}$ receives some constraining power from boosted top pair production, which in turn allows single top measurements to constrain $\mathcal{O}_{\phi Q}^{3}$.
@@
Altogether, we find that the public likelihoods do not have a significant effect on our SMEFT limits. As discussed in Sec. 3, the likelihoods available and included in our analysis all describe total cross sections, limiting their impact. On the positive side, public likelihoods allow for an accurate modelling of correlated systematics, an aspect we will discuss in Sec 4.2.
@@
Much of the improvement we see from our new dataset is due to the boosted regime. For SMEFT analyses such measurements are extremely helpful to constrain operators which include a momentum scaling. As discussed in Sec. 2.2, we add the unfolded ATLAS measurement of boosted top pair production [94]. In Fig. 1, we already showed the impact of a single SMEFT operator $\mathcal{O}_{Qd}^{8}$ on the normalized $m_{t\bar{t}}$ distribution of this measurement.
Here, we study its effect on the global analysis. Figure 10 demonstrates the impact on a selection of two-operator correlations. The complete analysis including all data in Tabs. 2, Tabs. 3 is compared to the case where the measurement of boosted tops from Ref. [94] is excluded. In the first panel, we observe an increase in constraining power on $\mathcal{O}_{Qq}^{18}$, while the constraints on $\mathcal{O}_{tG}$ are stable. This follows from the fact that this operator is instead constrained by the $t\bar{t}$ total cross section. In contrast, the limits on the Wilson coefficients for energy-growing 4-fermion operators improve by a factor two, as shown in the right hand panel for $\mathcal{O}_{Qq}^{8}$ and $\mathcal{O}_{td}^{8}$.
@@ 4.2 Correlated systematics
Public likelihoods, as discussed in Sec. 3, allow us to model and study correlated systematic uncertainties across measurements by the same experiment. For measurements without public likelihoods, the approximate treatment of correlations is discussed in Sec. 2.3. For the Higgs sector, we already know that the correlations of systematic uncertainties had a highly visible impact on the SMEFT analysis [65]. In particular, they lead to a marked shift in the most likely values of Wilson coefficients, while leaving the width of the limits unchanged.
@@
Here, we assess the impact of correlating systematic uncertainties on the top sector. In Fig. 11 we show two sets of constraints on a selection of Wilson coefficients. In blue, we show the constraints from a global analysis where all correlations between experimental systematics and between theory uncertainties are included. In black we show the same results, but treating all experimental systematics as uncorrelated. For all Wilson coefficients, we find good agreement, which indicates that in the top sector the correlations of systematic uncertainties cannot be ignored, but have limited effect on the final SMEFT limits.
We know that statistical uncertainties are not the leading challenge for global SMEFT analyses. So if the correlations between experimental systematics are not really relevant either, which uncertainties actually dominate the SMEFT analysis? While for the Higgs sector the modeling of theory uncertainties has surprisingly little effect on the SMEFT limits [65], the QCD nature of top pair production suggests that the situation will be different here. As a test, in Fig. 12 we repeat the comparison of Fig. 11, neglecting all theory uncertainties. As before, we show the global analysis with correlated systematics in blue, while in black these correlations are removed. Now we see a significant difference. When neglecting the correlations, we observe an increase in the size of the constraints as well as a sizeable shift in the most likely point. This is particularly marked in the 2-dimensional constraints on $C_{Qq}^{18}$ and $C_{tq}^{8}$.
Comparing Figs. 11 and 12 we learn the importance of theory uncertainties in the top sector. If we neglect the theory uncertainties, the effect of correlated systematics is non-negligible. While theory uncertainties currently wash out these effects, we expect them to become more important as SM calculations become more precise and theory uncertainties are reduced. Moreover, we cannot make any statement about the potential impact of public likelihoods for those kinematic measurements which drive the SMEFT sensitivity.
@@ 4.3 Marginalization vs profiling
As defined in Eq.(8), the central object of any SFitter analysis is a fully exclusive likelihood. It is evaluated over the combined space of Wilson coefficients and nuisance parameters. Obviously, the nuisance parameters are irrelevant for the physics interpretation of the global SMEFT analysis. In addition, we are usually not interested in showing all 22 Wilson coefficients at the same time, and instead reduce this space to one or two dimensions. Statistically, this can be done by profiling or marginalizing the likelihood. Only for perfect Gaussian distributions the two methods give the same results, as discussed in Sec. 2.3. In the Higgs-electroweak sector, significant deviations between the two methods appear through a large under-fluctuation in one bin of a kinematic distribution [65].
@@
Fig. 13 displays a selection of correlations, from a marginalization (black) and from profiling (blue) of the fully exclusive likelihood from all top sector measurements and Wilson coefficients. We show constraints for $\mathcal{O}_{tG}$ vs $\mathcal{O}_{tq}^{1}$ and $\mathcal{O}_{Qq}^{18}$ vs $\mathcal{O}_{Qd}^{8}$, but similar effects can be seen in many operator pairs. In general, marginalization leads to narrower constraints than profiling. This is particularly evident in the left panel of Fig. 13, and it is due to theory uncertainties and their flat likelihood distribution. With this choice, the profile likelihood can force a perfect agreement between data and predictions over a wide range of values for critical Wilson coefficients. When we marginalize over the exclusive likelihood, the difference between Gaussian and flat uncertainties is less pronounced, leading to more Gaussian and narrower one-dimensional distributions, as discussed in detail in Ref. [65]. This effect is especially visible in the top sector, where theory uncertainties are not only poorly defined [139], but also large.

### v3

@@ 3 Public likelihoods
Normalization (’normfactor’) $\kappa_{sb}=\mu$
Table 4: List of modifiers in the construction of the HistFactory likelihoods, adapted from Ref. [145]. Per-bin modifiers are denoted as $\gamma_{b}$, while interpolated modifiers are denoted as $\alpha$. Here $g_{p}$ and $f_{p}$ describe different interpolation strategies used to compute these from the values $\kappa_{sb,\alpha=\pm 1},\Delta_{sb,\alpha=\pm 1}$ provided in the likelihood. Luminosity and scale factors affect all bins equally and are denoted as $\lambda$ and $\mu$, respectively.
For a standard SFitter analysis, we extract systematic uncertainties for each measurement from the respective experimental publications. Systematics of the same type are fully correlated between measurements of the same experiment. This approach has drawbacks, for instance, we can only use the uncertainty categories reported in the experimental publications or on HEPData, and this information often needs to be extracted by hand. Public likelihoods include the full information on a large number of systematic uncertainties in a documented manner, making their implementation more accurate and efficient.
Likelihoods are published in the HistFactory format [56], similar to the SFitter likelihood in Eq.(8). For each bin $b$ measured in a kinematic distribution of a given channel or final state, it provides
@@
with the nominal expected rate $\bar{m}_{sb}$ and multiplicative $(\kappa_{sb})$ and additive $(\Delta_{sb})$ modifiers for each physics process $s$. These modifiers correspond to the nuisance parameters affecting the event rate $m_{b}$. The type of modifier and the constraints on its corresponding nuisance parameter depend on the type of uncertainty. The most common are given in Tab. 4. Using the public likelihoods in terms of modifiers and nominal rates $\bar{m}_{sb}$, we can reproduce the experimental results. For visualization, we use cabinetry [146], a python library making use of pyhf for statistical analyses.
Starting from the public likelihoods, we organize the full set of nuisance parameters corresponding to systematic uncertainties in a small number of categories. This allows for an easier numerical treatment at, essentially, no cost. To compute the ranges of nuisance parameters for these categories, we first use a profile likelihood to determine the central values and, in a second step, an analysis of the distribution of the nuisance parameter. In this section, we will show how we implement and test three public ATLAS likelihoods.
@@ 3.1 ATLAS $t\bar{t}$ likelihood
To test our implementation and evaluation of the public likelihood, we first reproduce some key ATLAS results in Fig. 2. We show the values for each nuisance parameter that maximizes the likelihood and the pulls,
@@
Here, $\hat{\theta}$ describes the maximum likelihood values and $\theta_{0}$ is the value before the fit, normalized to the pre-fit uncertainty $\Delta\theta$. We also show the impact of the individual nuisance parameters on the signal strength $\mu$. It is determined by repeating the fit after fixing the nuisance parameter to its maximum-likelihood value $\hat{\theta}$, shifted by its prefit (postfit) uncertainties ±$\pm\Delta\theta(\pm\Delta\hat{\theta})$. The left panel of Fig. 2 is taken from Ref. [47], while the right panel shows our reproduced results. Both sets show excellent agreement, with negligible differences for a few select nuisance parameters.
Next, we analyze the full likelihood as a function of a single nuisance parameter. This allows us to check the validity of a Gaussian likelihood, as assumed for systematic uncertainties in SFitter. For each nuisance parameter, we generally find excellent agreement with the Gaussian assumption, as shown on the left, with only a few exceptions. Figure 3 shows two such cases, one with excellent agreement and one with poor agreement. Even the larger deviations are under control, showing good agreement with the Gaussian approximation when we translate them into one standard deviation. Our combination of nuisance parameters into categories washes out non-Gaussian shapes in these exceptions.
Finally, we test the correlations between individual nuisance parameters and between nuisance parameters assigned to the categories implemented in SFitter. The left panel of Fig. 4 shows the correlations of all individual nuisance parameters with at least one correlation greater than 0.4. Since the public likelihoods do not provide additional metadata on all nuisance parameters, their labels do not necessarily match those used in the impact plots. We find that out of the many nuisance parameters included in the public likelihood, only very few are significantly correlated. We mainly see strong correlations, for instance, for modeling choices or jets.
In the standard SFitter approach, we group these individual nuisance parameters into uncorrelated categories and implement these categories with a single nuisance parameter each. Standard categories cover leptons, jets, tagging, and luminosity. Additional categories are process-specific, such as certain backgrounds or missing transverse energy. For all processes in our dataset, we use 21 nuisance parameters describing the systematic uncertainties assumed to be uncorrelated. Using the full likelihood, we show the correlations between these categories in the right panel of Fig. 4. The fact that the correlations between categories essentially vanish validates this SFitter approach.
@@ 3.2 ATLAS $t\bar{t}Z$ likelihood
The second likelihood we implement is for the $t\bar{t}Z$ rate measurement [48]. It simultaneously fits both 3-lepton and 4-lepton signal regions and the corresponding control regions. The parameter of interest is the $t\bar{t}Z$ signal strength. A total of 230 nuisance parameters describe the systematic uncertainties. Unlike for the $t\bar{t}$ likelihood, there are no uncertainties on the shape of the signal since each signal region is described by a single bin.
Following the method described for the $t\bar{t}$ analysis, we also test the $t\bar{t}Z$ likelihood and our implementation. Figure 5 compares the impact and pulls taken from Ref. [48] with those reproduced by us. We see excellent agreement for all nuisance parameters, with, at most, very minor differences.
As before, we then show the correlations between nuisance parameters with at least one correlation greater than 0.3 in the left panel of Fig. 6. We can compare them to our SFitter implementation on the right. We find that the correlations between individual nuisance parameters are already much smaller for this likelihood. The only strong correlations appear between scale uncertainties and the signal strength of the corresponding background. Consequently, the results, after combining all nuisance parameters into the SFitter categories, display negligible correlations between categories.
@@ 3.3 ATLAS $s$-channel single top likelihood
The correlations in the left panel of Fig. 8 show strong correlations only between select nuisance parameters. The strongest correlations appear between jet-related uncertainties and the signal strengths of the two dominant backgrounds, $t\bar{t}$ and $W$+jets. For SFitter, these nuisance parameters are put into the background uncertainty category. These strong correlations are therefore implicitly included in this larger category, and the final implementation into SFitter is essentially uncorrelated, as one can see in the right of Fig. 8. While one still finds a nonzero correlation between the jet and background uncertainties and between the jet and lepton uncertainties, these are all negligibly small.
Refer to caption Refer to caption
@@ 4 Global analysis
Using, for the first time, public likelihoods in a global SMEFT analysis allows us to look at different relevant questions. From the data included in our analysis, we know that our global analysis is somewhat unlikely to uncover a fundamental and statistically significant break-down of the SM. We first look at the impact on the constraining power from new measurements, especially boosted top kinematics, relative to Ref. [22]. We then study the impact of correlated uncertainties encoded in the public likelihoods. From a pure statistics perspective, we also check if lower-dimensional limits extracted by profiling and by marginalization differ. Finally, we provide SMEFT limits combining the updated top sector analysis with the electroweak and Higgs sector from Ref. [65].
@@ 4.1 Better and boosted measurements
Before we study more conceptual questions of global SMEFT analyses, we update our dataset with new measurements, as marked in Tabs. 2 and 3. In Fig. 9, we show the constraints on a selection of 2-dimensional correlations of Wilson coefficients, using all top data compared to the previous SFitter top analysis [22]. All constraints are the result of an analysis of all 22 Wilson coefficients. To extract limits on pairs of coefficients, we use a profile likelihood. Potential differences in marginalization will be discussed in Sec. 4.3.
@@
The right panel shows the improvement in constraints on $\mathcal{O}_{\phi Q}^{3}$ and $\mathcal{O}_{Qq}^{31}$. Single top production provides constraints on them, and we again benefit from the public likelihood [49], a new $p_{T,t}$ distribution in $t$-channel single top production [115], and new measurements of the $tW$ production cross section [121, 125]. We observe an improvement in the individual constraints and their correlation. In particular, $\mathcal{O}_{Qq}^{31}$ receives some constraining power from boosted top pair production, which in turn allows single top measurements to constrain $\mathcal{O}_{\phi Q}^{3}$.
@@
Altogether, we find that the public likelihoods do not have a significant effect on our SMEFT limits. As discussed in Sec. 3, the likelihoods available and included in our analysis all describe total cross sections, limiting their impact. On the positive side, public likelihoods allow for an accurate modeling of correlated systematics, an aspect we will discuss in Sec 4.2.
@@
Much of the improvement we see from our new dataset is due to the boosted regime. For SMEFT analyses, such measurements are extremely helpful to constrain operators which include momentum scaling. As discussed in Sec. 2.2, we add the unfolded ATLAS measurement of boosted top pair production [94]. In Fig. 1, we already showed the impact of a single SMEFT operator $\mathcal{O}_{Qd}^{8}$ on the normalized $m_{t\bar{t}}$ distribution of this measurement.
Here, we study its effect on the global analysis. Figure 10 demonstrates the impact on a selection of two-operator correlations. The complete analysis including all data in Tabs. 2, Tabs. 3 is compared to the case where the measurement of boosted tops from Ref. [94] is excluded. In the first panel, we observe an increase in constraining power on $\mathcal{O}_{Qq}^{18}$, while the constraints on $\mathcal{O}_{tG}$ are stable. This follows from the fact that this operator is instead constrained by the $t\bar{t}$ total cross section. In contrast, the limits on the Wilson coefficients for energy-growing 4-fermion operators improve by a factor of two, as shown in the right-hand panel for $\mathcal{O}_{Qq}^{8}$ and $\mathcal{O}_{td}^{8}$.
@@ 4.2 Correlated systematics
Public likelihoods, as discussed in Sec. 3, allow us to model and study correlated systematic uncertainties across measurements by the same experiment. For measurements without public likelihoods, the approximate treatment of correlations is discussed in Sec. 2.3. For the Higgs sector, we already know that the correlations of systematic uncertainties had a highly visible impact on the SMEFT analysis [65]. In particular, they lead to a marked shift in the most likely values of Wilson coefficients while leaving the width of the limits unchanged.
@@
Here, we assess the impact of correlating systematic uncertainties on the top sector. In Fig. 11, we show two sets of constraints on a selection of Wilson coefficients. In blue, we show the constraints from a global analysis where all correlations between experimental systematics and between theory uncertainties are included. In black we show the same results, but treating all experimental systematics as uncorrelated. For all Wilson coefficients, we find good agreement, which indicates that in the top sector, the correlations of systematic uncertainties cannot be ignored but have a limited effect on the final SMEFT limits.
We know that statistical uncertainties are not the leading challenge for global SMEFT analyses. So if the correlations between experimental systematics are not really relevant either, which uncertainties actually dominate the SMEFT analysis? While for the Higgs sector, the modeling of theory uncertainties has surprisingly little effect on the SMEFT limits [65], the QCD nature of top pair production suggests that the situation will be different here. As a test, in Fig. 12, we repeat the comparison of Fig. 11, neglecting all theory uncertainties. As before, we show the global analysis with correlated systematics in blue, while in black these correlations are removed. Now we see a significant difference. When neglecting the correlations, we observe an increase in the size of the constraints as well as a sizeable shift in the most likely point. This is particularly marked in the 2-dimensional constraints on $C_{Qq}^{18}$ and $C_{tq}^{8}$.
Comparing Figs. 11 and 12 we learn the importance of theory uncertainties in the top sector. If we neglect the theory uncertainties, the effect of correlated systematics is non-negligible. While theory uncertainties currently wash out these effects, we expect them to become more important as SM calculations become more precise, and theory uncertainties are reduced. Moreover, we cannot make any statement about the potential impact of public likelihoods for those kinematic measurements that drive the SMEFT sensitivity.
@@ 4.3 Marginalization vs profiling
As defined in Eq.(8), the central object of any SFitter analysis is a fully exclusive likelihood. It is evaluated over the combined space of Wilson coefficients and nuisance parameters. Obviously, the nuisance parameters are irrelevant to the physics interpretation of the global SMEFT analysis. In addition, we are usually not interested in showing all 22 Wilson coefficients at the same time and instead reduce this space to one or two dimensions. Statistically, this can be done by profiling or marginalizing the likelihood. Only for perfect Gaussian distributions do the two methods give the same results, as discussed in Sec. 2.3. In the Higgs-electroweak sector, significant deviations between the two methods appear through a large under-fluctuation in one bin of a kinematic distribution [65].
@@
Fig. 13 displays a selection of correlations from a marginalization (black) and profiling (blue) of the fully exclusive likelihood from all top sector measurements and Wilson coefficients. We show constraints for $\mathcal{O}_{tG}$ vs $\mathcal{O}_{tq}^{1}$ and $\mathcal{O}_{Qq}^{18}$ vs $\mathcal{O}_{Qd}^{8}$, but similar effects can be seen in many operator pairs. In general, marginalization leads to narrower constraints than profiling. This is particularly evident in the left panel of Fig. 13, and it is due to theory uncertainties and their flat likelihood distribution. With this choice, the profile likelihood can force a perfect agreement between data and predictions over a wide range of values for critical Wilson coefficients. When we marginalize over the exclusive likelihood, the difference between Gaussian and flat uncertainties is less pronounced, leading to more Gaussian and narrower one-dimensional distributions, as discussed in detail in Ref. [65]. This effect is especially visible in the top sector, where theory uncertainties are not only poorly defined [138], but also large.
4.4 Top-Higgs-electroweak combination
Figure 14: Results from a combined SMEFT analysis of the top sector and the Higgs-electroweak sector, showing the constraints at 95% CL on 43 degrees of freedom, resulting from a profiled likelihood.
Finally, making use of the numerical improvements in the SFitter implementation, we can combine the top-sector SMEFT analysis from this paper with the SFitter analysis of the Higgs, di-boson, and electroweak precision observables, Ref. [65]. This combination has been studied in the literature in detail, showing that the two sectors are linked, for instance, through $\mathcal{O}_{tG}$ [36, 37].
We confirm this state of the art and show the combined SFitter profile likelihood of the two sectors in Fig. 14. In total, 43 degrees of freedom are constrained: the 22 coefficients constrained by the top sector and discussed in Sec. 2.1, and 21 additional operators relevant to the Higgs, di-boson and electroweak observables. The notation and conventions for these 21 operators are provided in App. B. From the detailed discussion above and in Ref. [65], it is clear that the challenges and limitations of the global analyses in the two sectors are not the same. We show the limits at 95% CL from one-dimensional profile likelihoods of the combined fit (blue) and under the assumption of theory uncertainties reduced by a factor of 2 (orange). The numerical values of the constraints shown in Fig. 14 are provided in Tab. 6.
In the top sector, we find strong constraints on the four-fermion operators. The constraints on their Wilson coefficients are driven by kinematic distributions such as the ATLAS measurement of boosted top discussed in Sec. 2.2, and therefore theory uncertainties do not play an important role in their constraints. Conversely, the constraint on $C_{tG}$ improves significantly when theory uncertainties are halved, indicating that theory uncertainties dominate constraints obtained from top quark pair production total cross sections. Similarly, this hypothetical reduction of theory uncertainties has an effect on the constraints obtained from single top, $t\bar{t}W$ and $t\bar{t}Z$ on coefficients such $C_{tW}$, $C_{bW}$, and $C_{tZ}$.
On the other hand, we observe no significant changes in the constraints on the operators relevant to the Higgs, di-boson and electroweak sectors, shown in the lower half of Fig. 14, when theory uncertainties are reduced. The exception is $C_{\phi G}$, which also benefits from the top quark data through its correlation with $C_{tG}$ and $C_{t\phi}$. This is in agreement with Ref. [65], where it was found that in the Higgs-gauge sector, systematic uncertainties are the dominant source of uncertainty for many of the observables in this sector.